OpenStudy (anonymous):

determine c so that the function f(x)= 2cx+3 x<=1, and x^2+c x>1, is continous

jimthompson5910 (jim_thompson5910):

plug x = 1 into 2cx+3 to get 2c*1+3 = 2c+3 also, plug x = 1 into x^2+c to get 1^2+c = 1 + c Now equate those two parts: 2c+3 = 1+c and solve for c to get your answer

OpenStudy (anonymous):

\[ f(x) = \left\{ \begin{array}{lr} 2cx+3 & : x \leq 1\\ x^2+c& : x >1 \end{array} \right.\] just seeing if i could still do it

jimthompson5910 (jim_thompson5910):

you can also use \begin{cases} http://latex.wikia.com/wiki/Cases_%28LaTeX_environment%29 and you don't have to worry about having the "\right." in there (you still have to close off the cases though.

OpenStudy (anonymous):

Thanks! Got it!